OpenStudy (anonymous):

Derive \(y''+4y=cos \omega t\). Second order ODE problem.

OpenStudy (anonymous):

@amistre64 help please?

OpenStudy (anonymous):

Solve the homogeneous problem first (complex roots) and then make a guess for the function on the right hand side. I would try it with: \[\Large A\sin(\omega t) + B\cos ( \omega t) \]

OpenStudy (anonymous):

I just had a lecture on this but he was very unclear on this topic, basically he told us to replace \(y''+4y\) with an equation like \(Ly\) and then he lost me :(

OpenStudy (anonymous):

The LaPlace transform?

OpenStudy (amistre64):

there are many methods that can be used; so unless youre tied to a method ....

OpenStudy (anonymous):

sorry but the only thing I know was that we know \((Ly_p=g\) and \(L(c_1 y_1+c_2 y_2 )=0)\) and we want \(L(y_p+c_1 y_1+c_2 y_2 )=g\)

OpenStudy (amistre64):

im unsure about the method presented then. It does look laplacey, but its not ringing any bells for me.

OpenStudy (anonymous):

can you give me an example with the LaPlace transform? If there were any easier equations to solve with this method, I would gladly take it. I just want to understand this before we move on in class

OpenStudy (amistre64):

if it is laplace, then i fond pauls site to do a splendid job at being readable and concise. http://tutorial.math.lamar.edu/Classes/DE/LaplaceTransforms.aspx Personally, I would have gone with Limbuses suggestion for a solution.

OpenStudy (anonymous):

Thanks you! I will check that out.

OpenStudy (anonymous):

this is linear operator read more about linear operator. first find the complementary solution and general solution.

OpenStudy (anonymous):

solve the homogeneous equation

OpenStudy (anonymous):

and guess a general function that would give you a form of cosine when you take the first and the second derivative.

OpenStudy (amistre64):

hmmm, getusel is making more sense :) The "L" will later be used in LaPlace transforms, but at the moment it is most likely a generic operating notation.

OpenStudy (anonymous):

laplace is not represented by capital l for most cases rather by some distorted l

OpenStudy (anonymous):

ok so I am confused again. Do I solve \(y''+4y=0\) first?

OpenStudy (anonymous):

L is one of the linear operators like laplace.

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

go for the homogenous solution first y'' + 4y = 0 this will be the yc, or yh depending on textbook

OpenStudy (anonymous):

solve it

OpenStudy (anonymous):

using characterstics equation

OpenStudy (anonymous):

ok. \[\lambda ^{2} + 4\lambda =0\] so \(\lambda = 2i\) and \(y=c_{1}cos2t+c_{2}sin2t\) right?

OpenStudy (anonymous):

and then i can set \(c_1cos2t+c_2sin2t=cos \omega t\)?

OpenStudy (amistre64):

yes, but the way i was taught it was to consider that ca and c2 as functions; and get it down to the y'' again

OpenStudy (amistre64):

well, almost right; yp''+4yp = cos wt

OpenStudy (amistre64):

\[y_p=A(t)cos(2t)+B(t)sin(2t)\] \[y'_p=A'cos(2t)-A~2sin(2t)+B'sin(2t)+B~2cos(2t)\] but the A',B' terms equate to the homogenous part, and zero out leaveing us with \[y'_p=-A~2sin(2t)+B~2cos(2t)\] \[y''_p=-A'~2sin(2t)+B'~2cos(2t)-A~4cos(2t)-B~4sin(2t)\]

OpenStudy (amistre64):

now we have all the parts needed to express this in terms of y''p + 4yp= cos wt \[~~~~~y''_p=-A'~2sin(2t)+B'~2cos(2t)-A~4cos(2t)-B~4sin(2t)\\~+4y_p=~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~A~4cos(2t)+B~4sin(2t)\\---------------------------\\cos(wt)=-A'~2sin(2t)+B'~2cos(2t)\]

OpenStudy (amistre64):

what this does is give us 2 equations in 2 unknowns \[A'cos(2t)+B'sin(2t)=0\\-2A'sin(2t)+2B'cos(2t)=cos(wt)\] solve for A' adn B' and integrate them back into A and B

OpenStudy (anonymous):

This is very helpful. Thank you so much!

OpenStudy (amistre64):

theres a method called the wronskian what shortens this up alittle, since the repetitious stuff can be formulized into a matrix